OpenStudy (anonymous):

A window in a building is at a height of 10 m from the ground. The angle of depression of a point P on the ground from the window is 30 degree. The angle of elevation of the top of the building from the point P is 60 degrees. Find the height of the building.

OpenStudy (anonymous):

 from the figure, tan 30 = 10 / X Therefore, X = 10/tan30 And tan60 = H/X Therefore, H = Xtan60 = (10/tan30)*(tan60) = \[10*\sqrt{3}*\sqrt{3} = 10*3 = 30m \]